# Integrating Dynamical Systems Modeling with Spatiotemporal scRNA-seq Data Analysis

### v1

@@ 4.2.1 Discrete Spatiotemporal Dynamics Modeling
where $\mathbf{M}=\mathbf{D}^{\text{spa}}\odot\mathbf{D}^{\text{exp}}$ measures the distance between two cells in the same slice, and $\mathbf{D}^{\text{spa}}$ and $\mathbf{D}^{\text{exp}}$ are distance matrices constructed on each slice according to spatial coordinates $\mathbf{S}$ and gene expression $\mathbf{X}$, respectively. That is $d^{\text{spa}}_{i,j}=\|\mathbf{z}_{i}-\mathbf{z}_{j}\|_{2}$ and $d^{\text{exp}}_{i,j}=\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{2}$. In addition, the first and second terms in $\mathcal{E}^{\mathbf{M}}(\boldsymbol{\pi})$ promote the proximity of different descendants of a cell at the previous moment and the proximity of different ancestors of a cell at the later moment, respectively, and the third term is the usual GW term only replaces the spatial distance matrix $\mathbf{D}$ with the $\mathbf{M}$ distance matrix. In DeST-OT, the authors define the growth vector $\mathbf{\xi}=\boldsymbol{\pi}\mathbf{1}_{M}-\nu_{0}$ and the growth rate $\mathbf{g}=\log\left(1+N\mathbf{\xi}\right)/\left(t_{1}-t_{0}\right)$.
@@
Spateo (Qiu et al., 2022b) uses maps $\pi$ obtained by other OT-based methods to unify spatial coordinates at different times into the reference coordinate system by solving the generalized weighted Procrustes problem in Eq. (4.24) (possibly in 3D). Next, Spateo selects the spatial coordinates of the cell with the most weight at the late time point mapped from each cell at the early time point as its future state, formally
$\mathbf{z}_{i,\text{future}}=\mathbf{z}^{\prime}_{\arg\max\pi_{i,:}},$
@@
$\mathbf{v}^{\text{spa}}_{i}=\mathbf{z}_{i,\text{future}}-\mathbf{z}_{i}.$
Finally, Spateo recovers a continuous spatial velocity field $\frac{\mathrm{d}\mathbf{z}}{\mathrm{d}t}=f(\mathbf{z}),$ from the spatial velocity of each cell, allowing for a series of differential geometry analyses, including divergence, acceleration, curvature, and torsion.
@@ 4.2.2 Spatiotemporal Dynamics Modeling
The majority of current approaches model spatial coordinates based on Gromov-Wasserstein OT, which has no dynamic form. stVCR proposes to model spatial coordinates using rigid-body transformation invariant OT, as well as using the widely used Wasserstein OT for modeling gene expression and unbalanced OT for modeling cellular proliferation. Next, stVCR integrates all modules into dynamic forms, making it possible to reconstruct dynamic continuous trajectories of cell differentiation, migration and proliferation simultaneously.
@@
Rigid body transformation invariant optimal transport The method in (Cohen and Guibasm, 1999) considers the optimal transport problem invariant to a given set of manipulations $\mathcal{G}$. It simultaneously searches for the optimal mapping $\pi$ and the optimal transformation $g$ through the optimization problem:
@@
$(\boldsymbol{\pi}^{\star},g^{\star})=\mathop{\arg\min}_{\boldsymbol{\pi}\in\boldsymbol{\pi}(\nu_{0},\nu_{1}),g\in\mathcal{G}}\langle\mathbf{c}(g),\boldsymbol{\pi}\rangle\stackrel{{\scriptstyle\text{ def. }}}{{=}}\sum_{i,j}\pi_{i,j}d\big{(}\mathbf{z}_{i},g(\mathbf{Z}^{\prime}_{j})\big{)}.$ (4.25)
Solving problem (4.25) directly is difficult and it can be solved by iteratively
$\boldsymbol{\pi}^{(n)}$ $=\mathop{\arg\min}_{\boldsymbol{\pi}\in\boldsymbol{\pi}(\nu_{0},\nu_{1})}\sum_{i,j}\pi_{ij}d\big{(}\mathbf{z}_{i},g^{(n)}(\mathbf{Z}^{\prime}_{j})\big{)},$ (4.26)
$g^{(n+1)}$ $=\mathop{\arg\min}_{g\in\mathcal{G}}\sum_{i,j}\pi_{ij}^{(n)}d\big{(}\mathbf{z}_{i},g(\mathbf{Z}^{\prime}_{j})\big{)}.$ (4.27)
@@
The subproblem (4.26) is to solve a static OT. In addition, when we choose the set $\mathcal{G}$ as the set of rigid body transformations, we call this problem rigid body transformation invariant optimal transport. At this point, subproblem (4.27) is the generalized weighted Procrustes problem mentioned before.
@@
Consider the ST data $(\mathbf{Z}^{(0:K)},\mathbf{X}^{(0:K)})$ at $t_{0},t_{1}\dots t_{K}$ totaling $K$ time points, and the number of cells in each observation is $n_{0},n_{1}\dots n_{K}$. stVCR uses the spatial coordinate system of the data at $t_{0}$ as a reference, and searches for the optimal dynamics and the optimal rigid-body transformation $(\mathbf{r}_{1:k},R_{1:k})$ by interpolating the empirical probability distributions of the data after the rigid-body transformation $\hat{p}^{k}$ and the number of cells $n_{k}$ using a transport-with-growth partial differential equation (PDE)
$\partial_{t}p_{t}(\mathbf{z},\mathbf{x})+\nabla\cdot\Big{(}\big{(}\mathbf{v}_{t}(\mathbf{z},\mathbf{x}),\mathbf{b}_{t}(\mathbf{z},\mathbf{x})\big{)}p_{t}(\mathbf{z},\mathbf{x})\Big{)}=g_{t}(\mathbf{z},\mathbf{x})p_{t}(\mathbf{z},\mathbf{x}),$ (4.28)
@@
where $\mathbf{v}_{t}(\mathbf{z},\mathbf{x})$ is cell spatial migration velocity, $\mathbf{b}_{t}(\mathbf{z},\mathbf{x})$ is RNA velocity and $g_{t}(\mathbf{z},\mathbf{x})$ is cell growth rate. Thus the feasible state space $\mathcal{S}$ for the arguments under constraints is
$\mathcal{S}($ $\mathbf{Z}^{(0:K)},\mathbf{X}^{(0:K)}):=\big{\{}(p_{t},\mathbf{v}_{t},\mathbf{b}_{t},g_{t};\mathbf{R}_{1:K},\mathbf{r}_{1:K})\big{|}\ \partial_{t}p_{t}+\nabla\cdot((\mathbf{v}_{t},\mathbf{b}_{t})p_{t})=g_{t}p_{t},$ (4.29)
$p_{t_{0}}=p^{(0)},\|p_{t_{k}}\|_{1}=n_{k}/n_{0},\bar{p}_{t_{k}}={p}^{(k)},\mathbf{R}^{T}_{k}\mathbf{R}_{k}=\mathbf{I},\det\mathbf{R}_{k}=1,\ k=1,2,\ldots,K\big{\}},$
@@
where $\|p_{t}\|_{1}:=\int p_{t}\mathrm{d}\mathbf{z}\mathrm{d}\mathbf{x}$ is the total mass of $p_{t}$ and $\bar{p}_{t}:=p_{t}/\|p_{t}\|_{1}$. stVCR finds optimal dynamics $(p_{t},\mathbf{v}_{t},\mathbf{b}_{t},g_{t})$ and optimal rigid-body transformations $(\mathbf{R}_{1:K},\mathbf{r}_{1:K})$ by minimizing the Wasserstein-Fisher-Rao (WFR) distance
$\int_{t_{0}}^{t_{K}}\int_{\mathbb{R}^{G+d_{\text{spa}}}}\left(\left\|\mathbf{v}_{t}\right\|^{2}+\alpha_{\text{Exp}}\left\|\mathbf{b}_{t}\right\|^{2}+\alpha_{\text{Gro}}g^{2}_{t}\right)p_{t}(\mathbf{z},\mathbf{x})\mathrm{d}\mathbf{z}\mathrm{d}\mathbf{x}\mathrm{d}t$ (4.30)
@@
for $(p_{t},\mathbf{v}_{t},\mathbf{b}_{t},g_{t};\mathbf{R}_{1:K},\mathbf{r}_{1:K})\in\mathcal{S}(\mathbf{Z}^{(0:K)},\mathbf{X}^{(0:K)})$. According to the direct derivation of the solution of the Feynman-Kac type PDE (4.28) by characteristics, Eq. (4.30) has a dimensionally independent form
$\mathcal{L}_{\text{Dyn}}=\mathbb{E}_{(\mathbf{z}^{(t_{0})},\mathbf{x}^{(t_{0})})\sim p^{(0)}}\int_{t_{0}}^{t_{K}}\Big{(}$ $\|\mathbf{v}_{t}(\mathbf{z}^{(t)},\mathbf{x}^{(t)})\|^{2}+\alpha_{\text{Exp}}\|\mathbf{b}_{t}(\mathbf{z}^{(t)},\mathbf{x}^{(t)})\|^{2}$ (4.31)
$+\alpha_{\text{Gro}}\|g_{t}(\mathbf{z}^{(t)},\mathbf{x}^{(t)})\|^{2}\Big{)}w_{t}[\mathbf{z},\mathbf{x}]\mathrm{d}t,$
@@
where $\mathbf{z}^{(t)},\mathbf{x}^{(t)},w_{t}[\mathbf{z},\mathbf{x}]$ satisfies the characteristic ordinary differential equations (ODEs)
$\frac{d\mathbf{z}^{(t)}}{dt}=\mathbf{v}_{t}(\mathbf{z}^{(t)},\mathbf{x}^{(t)}),\ \frac{d\mathbf{x}^{(t)}}{dt}$ $=\mathbf{b}_{t}(\mathbf{z}^{(t)},\mathbf{x}^{(t)}),\ \frac{d\ln w_{t}}{dt}=g_{t}(\mathbf{z}^{(t)},\mathbf{x}^{(t)}),$ (4.32)
$(\mathbf{z}^{(t)},\mathbf{x}^{(t)},w_{t})|_{t=t_{0}}$ $=(\mathbf{z}^{(t_{0})},\mathbf{x}^{(t_{0})},1).$
@@
stVCR implemented the constraints $\|p_{t_{k}}\|_{1}=n_{k}/n_{0}$ and $\bar{p}_{t_{k}}=\hat{p}^{(k)}$ in Eq. (4.29) as soft penalties by performing distribution matching
$\mathcal{L}_{\text{Mch}}=\sum_{k=1}^{K}\big{(}W_{2}(\bar{p}_{t_{k}},\hat{p}^{(k)})\big{)}^{2}+\kappa_{\text{Gro}}\sum_{k=1}^{K}\frac{|\sum_{j=1}^{n_{0}}w_{t_{k},j}-n_{k}|}{n_{k}},$ (4.33)
@@
$c^{(k)}_{ij}=\kappa_{\text{Exp}}\|\mathbf{x}^{(t_{k})}_{i}-\hat{\mathbf{x}}^{(k)}_{j}\|_{2}^{2}+(1-\kappa_{\text{Exp}})\|\mathbf{z}^{(t_{k})}_{i}-\hat{\mathbf{z}}^{(k)}_{j}\|_{2}^{2},\quad i=1:n_{0},j=1:n_{k},$
@@
$\mathcal{L}=\mathcal{L}_{\text{Dyn}}+\lambda_{\text{Mch}}\mathcal{L}_{\text{Mch}}+\lambda_{\text{SSP}}\mathcal{L}_{\text{SSP}}^{\text{opt}}.$ (4.34)
In practice, stVCR parameterizes the dynamics $\mathbf{v}_{t}(\mathbf{z}^{(t)},\mathbf{x}^{(t)})$, $\mathbf{b}_{t}(\mathbf{z}^{(t)},\mathbf{x}^{(t)})$ and $g_{t}(\mathbf{z}^{(t)},\mathbf{x}^{(t)})$ into neural networks as well as parameterizes the rotation matrix $\mathbf{R}_{1:K}$ into rotation angles $\alpha_{1:K}$ (or Euler angles $(\alpha_{1:K},\beta_{1:K},\gamma_{1:K})$ in 3D) and solves iteratively using back-propagation algorithm.
5 Extensions, Challenges and Future Directions

### v2

@@ 4.2.1 Discrete Spatiotemporal Dynamics Modeling
where $\boldsymbol{M}=\boldsymbol{D}^{\text{spa}}\odot\boldsymbol{D}^{\text{exp}}$ measures the distance between two cells in the same slice, and $\boldsymbol{D}^{\text{spa}}$ and $\boldsymbol{D}^{\text{exp}}$ are distance matrices constructed on each slice according to spatial coordinates $\boldsymbol{S}$ and gene expression $\boldsymbol{X}$, respectively. That is, $d^{\text{spa}}_{i,j}=\|\boldsymbol{z}_{i}-\boldsymbol{z}_{j}\|_{2}$ and $d^{\text{exp}}_{i,j}=\|\boldsymbol{x}_{i}-\boldsymbol{x}_{j}\|_{2}$. In addition, the first and second terms in $\mathcal{E}^{\boldsymbol{M}}(\boldsymbol{\pi})$ promote the proximity of different descendants of a cell at the previous moment and the proximity of different ancestors of a cell at the later moment, respectively, and the third term is the usual GW term that only replaces the spatial distance matrix $\boldsymbol{D}$ with the $\boldsymbol{M}$ distance matrix. In DeST-OT, the authors define the growth vector $\boldsymbol{\xi}=\boldsymbol{\pi}\boldsymbol{1}_{M}-\nu_{0}$ and the growth rate $\boldsymbol{g}=\log\left(1+N\boldsymbol{\xi}\right)/\left(t_{1}-t_{0}\right)$.
@@
Spateo Qiu et al. (2022) uses maps $\pi$ obtained by other OT-based methods to unify spatial coordinates at different times into the reference coordinate system by solving the generalized weighted Procrustes problem in Equation (25) (possibly in 3D). Next, Spateo selects the spatial coordinates of the cell with the most weight at the late time point mapped from each cell at the early time point as its future state, formally
$\boldsymbol{z}_{i,\text{future}}=\boldsymbol{z}^{\prime}_{\arg\max\pi_{i,:}},$
@@
$\boldsymbol{v}^{\text{spa}}_{i}=\boldsymbol{z}_{i,\text{future}}-\boldsymbol{z}_{i}.$
Finally, Spateo recovers a continuous spatial velocity field $\frac{\mathrm{d}\boldsymbol{z}}{\mathrm{d}t}=f(\boldsymbol{z}),$ from the spatial velocity of each cell, allowing for a series of differential geometry analyses, including divergence, acceleration, curvature, and torsion.
@@ 4.2.2 Spatiotemporal Dynamics Modeling
The majority of current approaches model spatial coordinates based on Gromov–Wasserstein OT, which has no dynamic form. Recently, stVCR Peng et al. (2024) proposes to model spatial coordinates using rigid-body transformation invariant OT, as well as using the widely used Wasserstein OT for modeling gene expression and unbalanced OT for modeling cellular proliferation. Next, stVCR integrates all modules into dynamic forms, making it possible to reconstruct dynamic continuous trajectories of cell differentiation, migration, and proliferation simultaneously.
@@
Rigid body transformation invariant optimal transport The method in Cohen and Guibasm (1999) considers the optimal transport problem invariant to a given set of manipulations $\mathcal{G}$. It simultaneously searches for the optimal mapping $\pi$ and the optimal transformation $g$ through the optimization problem:
@@
$(\boldsymbol{\pi}^{\star},g^{\star})=\mathop{\arg\min}_{\boldsymbol{\pi}\in\boldsymbol{\Pi}(\nu_{0},\nu_{1}),g\in\mathcal{G}}\langle\boldsymbol{c}(g),\boldsymbol{\pi}\rangle\stackrel{{\scriptstyle\text{ def. }}}{{=}}\sum_{i,j}\pi_{i,j}d\big{(}\boldsymbol{z}_{i},g(\boldsymbol{Z}^{\prime}_{j})\big{)}.$ (26)
Solving problem (26) directly is difficult, and it can be solved iteratively by
$\boldsymbol{\pi}^{(n)}$ $=\mathop{\arg\min}_{\boldsymbol{\pi}\in\boldsymbol{\Pi}(\nu_{0},\nu_{1})}\sum_{i,j}\pi_{ij}d\big{(}\boldsymbol{z}_{i},g^{(n)}(\boldsymbol{Z}^{\prime}_{j})\big{)},$ (27)
$g^{(n+1)}$ $=\mathop{\arg\min}_{g\in\mathcal{G}}\sum_{i,j}\pi_{ij}^{(n)}d\big{(}\boldsymbol{z}_{i},g(\boldsymbol{Z}^{\prime}_{j})\big{)}.$ (28)
@@
The subproblem (27) is to solve a static OT. In addition, when we choose the set $\mathcal{G}$ as the set of rigid body transformations, we call this problem rigid body transformation invariant optimal transport. At this point, subproblem (28) is the generalized weighted Procrustes problem mentioned before.
@@
Consider the ST data $(\boldsymbol{Z}^{(0:K)},\boldsymbol{X}^{(0:K)})$ at $t_{0},t_{1}\dots t_{K}$ totaling $K$ time points, and the number of cells in each observation is $n_{0},n_{1}\dots n_{K}$. stVCR uses the spatial coordinate system of the data at $t_{0}$ as a reference, and searches for the optimal dynamics and the optimal rigid-body transformation $(\boldsymbol{r}_{1:k},R_{1:k})$ by interpolating the empirical probability distributions of the data after the rigid-body transformation $\hat{p}^{k}$ and the number of cells $n_{k}$ using a transport-with-growth partial differential equation (PDE)
$\partial_{t}p_{t}(\boldsymbol{z},\boldsymbol{x})+\nabla\cdot\Big{(}\big{(}\boldsymbol{v}_{t}(\boldsymbol{z},\boldsymbol{x}),\boldsymbol{b}_{t}(\boldsymbol{z},\boldsymbol{x})\big{)}p_{t}(\boldsymbol{z},\boldsymbol{x})\Big{)}=g_{t}(\boldsymbol{z},\boldsymbol{x})p_{t}(\boldsymbol{z},\boldsymbol{x}),$ (29)
@@
where $\boldsymbol{v}_{t}(\boldsymbol{z},\boldsymbol{x})$ is cell spatial migration velocity, $\boldsymbol{b}_{t}(\boldsymbol{z},\boldsymbol{x})$ is RNA velocity and $g_{t}(\boldsymbol{z},\boldsymbol{x})$ is cell growth rate. Thus, the feasible state space $\mathcal{S}$ for the arguments under constraints is
$\mathcal{S}($ $\boldsymbol{Z}^{(0:K)},\boldsymbol{X}^{(0:K)}):=\big{\{}(p_{t},\boldsymbol{v}_{t},\boldsymbol{b}_{t},g_{t};\boldsymbol{R}_{1:K},\boldsymbol{r}_{1:K})\big{|}\ \partial_{t}p_{t}+\nabla\cdot((\boldsymbol{v}_{t},\boldsymbol{b}_{t})p_{t})=g_{t}p_{t},$ (30)
$p_{t_{0}}=p^{(0)},\|p_{t_{k}}\|_{1}=n_{k}/n_{0},\bar{p}_{t_{k}}={p}^{(k)},\boldsymbol{R}^{T}_{k}\boldsymbol{R}_{k}=\boldsymbol{I},\det\boldsymbol{R}_{k}=1,\ k=1,2,\ldots,K\big{\}},$
@@
where $\|p_{t}\|_{1}:=\int p_{t}\mathrm{d}\boldsymbol{z}\mathrm{d}\boldsymbol{x}$ is the total mass of $p_{t}$ and $\bar{p}_{t}:=p_{t}/\|p_{t}\|_{1}$. stVCR finds optimal dynamics $(p_{t},\boldsymbol{v}_{t},\boldsymbol{b}_{t},g_{t})$ and optimal rigid-body transformations $(\boldsymbol{R}_{1:K},\boldsymbol{r}_{1:K})$ by minimizing the Wasserstein–Fisher–Rao (WFR) distance
$\int_{t_{0}}^{t_{K}}\int_{\mathbb{R}^{G+d_{\text{spa}}}}\left(\left\|\boldsymbol{v}_{t}\right\|^{2}+\alpha_{\text{Exp}}\left\|\boldsymbol{b}_{t}\right\|^{2}+\alpha_{\text{Gro}}g^{2}_{t}\right)p_{t}(\boldsymbol{z},\boldsymbol{x})\mathrm{d}\boldsymbol{z}\mathrm{d}\boldsymbol{x}\mathrm{d}t$ (31)
@@
for $(p_{t},\boldsymbol{v}_{t},\boldsymbol{b}_{t},g_{t};\boldsymbol{R}_{1:K},\boldsymbol{r}_{1:K})\in\mathcal{S}(\boldsymbol{Z}^{(0:K)},\boldsymbol{X}^{(0:K)})$. According to the direct derivation of the solution of the Feynman–Kac type PDE (29) by characteristics, Equation (31) has a dimensionally independent form
$\mathcal{L}_{\text{Dyn}}=\mathbb{E}_{(\boldsymbol{z}^{(t_{0})},\boldsymbol{x}^{(t_{0})})\sim p^{(0)}}\int_{t_{0}}^{t_{K}}\Big{(}$ $\|\boldsymbol{v}_{t}(\boldsymbol{z}^{(t)},\boldsymbol{x}^{(t)})\|^{2}+\alpha_{\text{Exp}}\|\boldsymbol{b}_{t}(\boldsymbol{z}^{(t)},\boldsymbol{x}^{(t)})\|^{2}$ (32)
$+\alpha_{\text{Gro}}\|g_{t}(\boldsymbol{z}^{(t)},\boldsymbol{x}^{(t)})\|^{2}\Big{)}w_{t}[\boldsymbol{z},\boldsymbol{x}]\mathrm{d}t,$
@@
where $\boldsymbol{z}^{(t)},\boldsymbol{x}^{(t)},w_{t}[\boldsymbol{z},\boldsymbol{x}]$ satisfies the characteristic ordinary differential equations (ODEs)
$\frac{d\boldsymbol{z}^{(t)}}{dt}=\boldsymbol{v}_{t}(\boldsymbol{z}^{(t)},\boldsymbol{x}^{(t)}),\ \frac{d\boldsymbol{x}^{(t)}}{dt}$ $=\boldsymbol{b}_{t}(\boldsymbol{z}^{(t)},\boldsymbol{x}^{(t)}),\ \frac{d\ln w_{t}}{dt}=g_{t}(\boldsymbol{z}^{(t)},\boldsymbol{x}^{(t)}),$ (33)
$(\boldsymbol{z}^{(t)},\boldsymbol{x}^{(t)},w_{t})|_{t=t_{0}}$ $=(\boldsymbol{z}^{(t_{0})},\boldsymbol{x}^{(t_{0})},1).$
@@
stVCR implemented the constraints $\|p_{t_{k}}\|_{1}=n_{k}/n_{0}$ and $\bar{p}_{t_{k}}=\hat{p}^{(k)}$ in Equation (30) as soft penalties by performing distribution matching
$\mathcal{L}_{\text{Mch}}=\sum_{k=1}^{K}\big{(}W_{2}(\bar{p}_{t_{k}},\hat{p}^{(k)})\big{)}^{2}+\kappa_{\text{Gro}}\sum_{k=1}^{K}\frac{|\sum_{j=1}^{n_{0}}w_{t_{k},j}-n_{k}|}{n_{k}},$ (34)
@@
$c^{(k)}_{ij}=\kappa_{\text{Exp}}\|\boldsymbol{x}^{(t_{k})}_{i}-\hat{\boldsymbol{x}}^{(k)}_{j}\|_{2}^{2}+(1-\kappa_{\text{Exp}})\|\boldsymbol{z}^{(t_{k})}_{i}-\hat{\boldsymbol{z}}^{(k)}_{j}\|_{2}^{2},\quad i=1:n_{0},j=1:n_{k},$
@@
$\mathcal{L}=\mathcal{L}_{\text{Dyn}}+\lambda_{\text{Mch}}\mathcal{L}_{\text{Mch}}+\lambda_{\text{SSP}}\mathcal{L}_{\text{SSP}}^{\text{opt}}.$ (35)
In practice, stVCR parameterizes the dynamics $\boldsymbol{v}_{t}(\boldsymbol{z}^{(t)},\boldsymbol{x}^{(t)})$, $\boldsymbol{b}_{t}(\boldsymbol{z}^{(t)},\boldsymbol{x}^{(t)})$ and $g_{t}(\boldsymbol{z}^{(t)},\boldsymbol{x}^{(t)})$ into neural networks as well as parameterizes the rotation matrix $\boldsymbol{R}_{1:K}$ into rotation angles $\alpha_{1:K}$ (or Euler angles $(\alpha_{1:K},\beta_{1:K},\gamma_{1:K})$ in 3D) and solves iteratively using back-propagation algorithm.
5 Extensions, Challenges, and Future Directions